A block of mass m1 rests on a horizontal table. A light string connected to it passes over a frictionless pulley at the edge of table and from its other end another block B of mass m2 is suspended.The coefficient of Kinetic friction between block and the table is given.When the block A is sliding on the table,the tension in the string is given by _____________.
ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย (AIPMT-2015)

Grade:11

1 Answers

Arun
5 years ago
Dear Student ,
here in this question the m1ย rests on the table and the mass m2ย is connected by a light string .
Let the tension on the string is T .
So , when the block A starts to slide then ,
T - F' = m1a ....(1)
and m2g -T = m2a ....(2)
Now , F' = frictional force = Mu * m * g
km1g
Now solving equations 1 and 2 we get
ย 
a = m2g โ€“ mu * m1* g / m1 + m2
ย 
put this value in 2
ย 
You will get the required tensuion